OpenStudy (anonymous):

1. Calculate the average atomic mass of rubidium. Rubidium has two isotopes, 85Rb and 87Rb. 85Rb has an atomic mass of 84.912 amu and occurs at an abundance of 72.17%. 87Rb has an atomic mass of 86.909 amu and occurs at an abundance of 27.83%. Show your work.

OpenStudy (aaronq):

you should be asking this in the chem section, but use this formula: average atomic mass=\(\Sigma(mass_{Isotope A}*\%abundance_{Isotope A}) \)

OpenStudy (anonymous):

Is it possible that you write that more clearly?

OpenStudy (aaronq):

"\(\Sigma\)" means summation (addition) you're basically taking the weighted average of the isotopes. so multiply each isotope's mass by it's abundance and add the values that you get.

OpenStudy (anonymous):

72.1752 + 75.61083 = 150.78603 So would 150.78603 be the average atomic mass of rubidium?

OpenStudy (aaronq):

hmm that doesn't seem right, are you sure you multiplied each mass by it's own abundance?

OpenStudy (anonymous):

Yes

OpenStudy (aaronq):

i don't think you did it right, it should be: 0.7217*84.912+86.909*0.2783=85.4677651

OpenStudy (anonymous):

Do you do anything with 85Rb, and 87Rb?

OpenStudy (aaronq):

those are the names of the isotopes (e.g. \(^{85}Rb\) for isotope with 84.912 amu mass), their masses and their abundance is what you need to use.

OpenStudy (anonymous):

How does this look? Average Atomic Mass = (Mass of isotope A ∗ % Abundance of isotope A) + (Mass of isotope B ∗ % Abundance of isotope B) Average Atomic Mass = (84.912 ∗ 0.7217) + (86.909 ∗ 0.2783) Average Atomic Mass = 61.2809904 + 24.1867747 85.4677651 = 61.2809904 + 24.1867747 Average Atomic Mass = 85.4677651

OpenStudy (aaronq):

thats right, dude. \(\checkmark\)

OpenStudy (anonymous):

Thanks soo much for your help!

OpenStudy (aaronq):

no problem !